# Proximal Interacting Particle Langevin Algorithms

### v2

@@ 5.3.2 Comparing proximal algorithms with IPLA
MYIPLA $2.00$ $-0.107$ $22$ $2.27$
MYPGD $\mathbf{1.50}$ $-0.102$ $20$ $2.02$
PIPGLA $2.00$ $\mathbf{-0.096}$ $33$ $\mathbf{1.73}$
PGD $2.00$ $-0.098$ $\mathbf{19}$ $8.80$
IPLA $1.99$ $-0.101$ $\mathbf{19}$ $11.70$
5.3.3 Overall performances
We analyse the test errors and log pointwise predictive density (LPD) achieved using the final particle cloud $X_{500}^{1},\dots,X_{500}^{N}$ for MYIPLA, MYPGD and PIPGLA, with $N=1,10,100$, and corresponding computation times averaged over 10 replicates.
The test errors (Table 5) have similar values to that of PGD under the setting of Kuntz et al. [2023, Table 2], which is similar to ours but with Normal instead of Laplace priors on the weights. We observe in Table 5 that performance improves as the number of particles $N$ increases, especially for PIPGLA. When the number of particles is $100$, the three algorithms MYIPLA, PIPGLA and MYPGD have similar performance and runtimes, with MYIPLA having a slightly better performance and PIPGLA being slightly faster.
Table 5: Bayesian neural network. Test errors achieved using the final particle cloud $X_{500}^{1},\dots,X_{500}^{N}$ for MYIPLA, MYPGD and PIPGLA, with $N=1,10,100$, and corresponding computation times averaged over 10 replicates. The hightlighted value show the best error and runtime in the case $N=100$.
[…]
MYIPLA 1 $7.90\pm 5.17$ $1.84\pm 0.03$
10 $3.35\pm 1.05$ $10.43\pm 0.48$
100 $\mathbf{2.45\pm 0.91}$ $89.72\pm 3.37$
PIPGLA 1 $21.65\pm 9.88$ $1.83\pm 0.04$
10 $7.45\pm 2.13$ $9.59\pm 0.15$
100 $2.65\pm 1.32$ $\mathbf{81.98\pm 1.64}$
MYPGD 1 $5.15\pm 1.48$ $1.90\pm 0.06$
10 $3.15\pm 1.14$ $10.40\pm 0.46$
100 $2.55\pm 0.90$ $83.20\pm 6.01$
5.4 Sparse matrix completion
In this section, we demonstrate another application of our methods: the problem of matrix completion. Matrix completion [Candès and Plan, 2010; Liu et al., 2018] focuses on recovering an intact matrix with low-rank property from incomplete data. Its application varies from wireless communications [Kortas et al., 2017], traffic sensing [Mardani and Giannakis, 2014] to integrated radar and recommender systems [Gogna and Majumdar, 2015]. We consider two different settings for this problem and analyse the performance of MYIPLA.
5.4.1 Example 1
Given matrices $Y^{l}=\Phi X^{l}+\varepsilon^{l}\in\mathbb{R}^{m\times n}$, where $X^{l}\in\mathbb{R}^{r\times n}$ is sparse, $\Phi\in\mathbb{R}^{m\times r}$, $r\ll m<n$ and $\varepsilon^{l}$ has normally distributed entries $\varepsilon_{i,j}^{l}\sim\mathcal{N}(0,\sigma^{2})$ with known variance $\sigma^{2}$. Let $\Omega^{l}$ denote the set of observed entries for matrix $Y^{l}$, that is, suppose we only observe $Y_{ij}^{l}$, $(i,j)\in\Omega^{l},l=1,\dots,L$. Our goal is to estimate the missing entries. This problem can be restated using the following probability density
$p_{\beta}(Y,X)=p_{\beta}(X)p_{\beta}(Y|X)=C(\theta)e^{-e^{-\theta}\sum_{l}\|X^{l}\|_{\text{tr}}}\prod_{l}\prod_{(i,j)\in\Omega^{l}}\mathcal{N}(Y_{ij}|\Phi X_{ij}^{l},\sigma^{2}),$
where $\beta=(\theta,\Phi)$, $p_{\beta}(X)$ is a sparsity-inducing prior and $\|\cdot\|_{\text{tr}}$ is the trace (or nuclear) norm, which is a convex envelope of the rank function [Bach, 2008], and is defined as
$\|X\|_{\text{tr}}=\sum_{i=1}^{r}\sigma_{i}(X),$
$r=\text{rank}(X)$ and $\sigma_{1}(X)\geq\dots\geq\sigma_{r}(X)\geq 0$ are the singular values. The resulting prior can be seen as a regularisation term [Grave et al., 2011; Boursier et al., 2022]. The negative log density or potential is decomposed as
$U(\theta,\Phi,X)=\underbrace{-\log(C(\theta))+e^{-\theta}\sum_{l}\|X^{l}\|_{\text{tr}}}_{g_{2}(\beta,X)}+\underbrace{\sum_{l}\sum_{(i,j)\in\Omega^{l}}\frac{(Y_{ij}-\Phi X_{ij}^{l})^{2}}{2}}_{g_{1}(\beta,X)}.$
To implement MYIPLA we need to compute the proximal operator of $g_{2}$ which can be found in Appendix D.4.1. The data used for this example is described in the same appendix.
To analyse the performance of the algorithm, we compute the normalised MSE for the missing entries, that is,
$\frac{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}-P_{\Omega^{\mathsf{c}}}Y_{K}^{N}\|^{2}}{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}\|^{2}},$
where $P_{\Omega^{\mathsf{c}}}$ denotes the projection operator over the unobserved entries. The normalised error of the missing entries for the final particle cloud is $3.11\times 10^{-2}$, where we have considered $N=30$ particles and $K=500$ steps. The last value of the $\theta$ iterates is $\theta=0.1$, which provides a data-driven estimation of the regularisation parameter.
5.4.2 Example 2
Let $Y$ be the following matrix $Y=\Phi X+\varepsilon$, where $\;Y\in\mathbb{R}^{m\times n},\;X\in\mathbb{R}^{r\times n}$ is sparse, $\Phi\in\mathbb{R}^{m\times r}$ with $r\ll m<n$ and $\varepsilon$ has normally distributed entries $\varepsilon_{i,j}\sim\mathcal{N}(0,\sigma^{2})$ with known variance $\sigma^{2}$. Suppose we only observe entries $Y_{ij}$, $(i,j)\in\Omega$, where $\Omega$ denote the set of indices for the observed entries. We want to recover the matrix $\Phi X$ and also estimate the distribution of the entries of $\Phi$, which we assume follow a Normal distribution with mean $\theta$ and variance $\sigma^{2}$, that is, $\Phi_{i,j}\sim\mathcal{N}(\theta,\tilde{\sigma}^{2})$, where $\tilde{\sigma}$ is known. This problem can be formulated using the following probability density
$p_{\beta}(Y,X,\Phi)$ $=p_{\beta}(X)p_{\beta}(\Phi)p_{\beta}(Y|X,\Phi)$
$=C(\xi)e^{-e^{-\xi}\|X\|_{\text{tr}}}\prod_{i,j}\mathcal{N}(\Phi_{ij}|\theta,\tilde{\sigma}^{2})\prod_{(i,j)\in\Omega}\mathcal{N}(Y_{ij}|(\Phi X)_{ij},{\sigma}^{2}),$
where $\beta=(\xi,\theta)$ and $p_{\beta}(X)$ is a sparsity-inducing prior. The negative log density or potential can be expressed as
$U(\beta,X,\Phi)=\underbrace{-\log(C(\xi))+e^{-\xi}\|X\|_{\text{tr}}}_{g_{2}(\beta,X,\Phi)}+\underbrace{\sum_{i,j}\frac{(\Phi_{ij}-\theta)^{2}}{2\tilde{\sigma}^{2}}+\sum_{(i,j)\in\Omega}\frac{(Y_{ij}-(\Phi X)_{ij})^{2}}{2{\sigma}^{2}}}_{g_{1}(\beta,X,\Phi)}.$
To estimate $\beta$ from the data, we compute the proximal operator of $g_{2}$ in Appendix D.4.2. The data generation process is also described in this appendix. It is important to highlight that in this example we consider a cloud of particles for both $X$ and $\Phi$.
To evaluate the performance, we compute the normalised MSE for the missing entries of $Y$, the average normalised MSE of the final particle cloud $\Phi_{K}^{1},\dots,\Phi_{K}^{N}$, as well as the normalised MSE of a estimator of $\theta$ given by the last value of the iterates of the $\theta$ chain (Table 6). The evolution of the performance metrics across steps is shown in Figure 9, Appendix D.4.2.
Table 6: Performance metrics for the sparse matrix completion Example 2, when the true value of $\theta=1.4$ for the final particle cloud for $N=30$ particles and $K=500$ steps.
[…]
4.70 8.94 9.69
6 Conclusions
This work focuses on the maximum marginal likelihood problem for non-differentiable latent variable models arising in many machine learning and signal processing applications. Our main contribution is to propose an interacting particle Langevin algorithm, extending the methods proposed for differentiable latent variable models [Kuntz et al., 2023; Akyildiz et al., 2023]. To do so, we have leveraged proximal methods [Parikh and Boyd, 2014; Combettes and Pesquet, 2011], akin to proximal MCMC methods [Pereyra, 2016]. We have analysed the resulting family of algorithms and provided nonasymptotic analysis bounding the optimisation error $\mathbb{E}[\|\theta_{n}^{N}-\theta^{\star}\|^{2}]^{1/2}$. These results are direct analogues of sampling methods within the non-differentiable setting, extended to the case of estimation in latent variable models where the problem is a mixture of sampling and optimisation.
We have demonstrated the applicability and efficiency of our methods on different numerical examples. Specifically, we have conducted four experiments, the first involving a hierarchical toy model in which all assumptions are satisfied to verify the performance of our methods. Secondly, a Bayesian logistic regression model, in which we have studied the differences between using an iterative approach or an approximation of $\text{prox}_{g_{2}}^{\lambda}$, concluding that although their accuracy is very similar, the approximate methods are considerably faster, specifically by around 25%. Next, we have considered a classification problem using a Bayesian neural network, whose posteriors are highly multimodal, with sparsity-inducing priors on the weights. We have analysed the advantages and performance of using Laplace priors versus Normal priors, and investigated how pruning affects performance in both cases. We have also compared how proximal algorithms lead to more concentrated posteriors compared to algorithms which ignore the set of non-differentiable points. Finally, we have dealt with sparse matrix completion problems, which has wide ranging applications in machine learning.
We note that, for the cases where the set of non-differentiable points is measure zero, our methods outperform the standard methods which ignores this and uses standard gradient based methods. In particular, these methods, adapted to our setting, would have similar theoretical guarantees [Johnston and Sabanis, 2023], however, the use of proximal operators and proximal methods significantly outperform subgradient based variants. In particular, using sparsity-inducing priors (like Laplace priors) the posterior distribution is more concentrated when employing proximal interacting particle Langevin algorithms compared to Langevin algorithms such as those in Akyildiz et al. [2023]; Kuntz et al. [2023], as we showed in the Bayesian neural network experiment (Section 5.3).
Our algorithms present a novel approach to work with latent variable models with sparsity-inducing priors, priors with compact support such as uniform priors, among others. This allows working with Bayesian models arising from different types of non-differentiable regularisations, such as the Lasso regularisation, the elastic net or the total-variation norm, to name a few.
Future work holds many promising avenues. First, we envisage our methods to be expanded beyond the convex setting (as they already perform well in the non-convex empirical examples as we have shown in Section 5.3). Our theoretical framework can be extended to the non-convex case, using recent non-convex optimisation bounds [Zhang et al., 2023; Akyildiz and Sabanis, 2024] and their non-differentiable adaptations. We can further adapt stochastic gradients straightforwardly as in Akyildiz et al. [2023]. This is especially important in a machine learning context, as gradients are calculated using a mini-batch of data to improve efficiency. Our novel bounds on the difference between the true minimiser and the minimiser of the Moreau-Yosida approximation can also be used within the recent multiscale approaches to extend them to non-differentiable settings, see, e.g., Akyildiz et al. [2024].
Acknowledgments and Disclosure of Funding
@@ 6 Conclusions
PCE is supported by EPSRC through the Modern Statistics and Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1.
@@ A.1 Approximation of minimisers
Before proceeding with the proof of Theorem 4.1 and 4.2, we derive a result controlling the distance between the maximiser of $p_{\theta}(y)=\int_{{}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\star}$, and the maximiser of $p_{\theta}^{\lambda}(y)=\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\lambda,\star}$.
@@
$k_{\lambda}(\theta):=\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x,\qquad\qquad k(\theta):=\int_{{}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$
@@
i.e. positive definite due to Remark 4.2 and assumption A5. Thus the Jacobian of $\phi$ w.r.t. $w$ is invertible. Hence, the implicit function theorem shows that there is some locally defined $\zeta$, such that $\zeta(0)=v$ and $\phi(\lambda,\zeta(\lambda))=0$. Furthermore,
@@ A.2 MYIPLA
The proof follows from Karatzas and Shreve [1991] and Akyildiz et al. [2023, Proposition 1]. ∎
@@ A.3 PIPULA
for every $x,y\in\mathbb{R}^{d}$. By plugging this result into the proof of Akyildiz et al. [2023, Lemma B.1] we obtain an equivalent result to that of Akyildiz et al. [2023, Proposition 5]:
@@ A.4.1 Error bound for proximal gradient Langevin algorithm
For convenience we consider the following decomposition of the PGLA update targeting a distribution $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}^{\lambda}))$ over d
@@
We are going to derive a bound for $W_{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})$, where $\mathcal{L}({X_{n+1}})$ denotes the distribution of the random variable $X_{n+1}$. For every $\pi$-integrable function $g:\mathbb{R}^{d}\to\mathbb{R}$, we define $\mathcal{E}_{g}(\pi)=\int g\mathrm{d}\pi$ and we denote $\mathcal{F}=\mathcal{E}_{g_{1}+g_{2}}+\mathcal{H}$, where $\mathcal{H}$ is the negative entropy $\mathcal{H}(\pi)=\int\log\pi\;\mathrm{d}\pi$. We also introduce the subdifferential of a convex function and its minimal section, since we will use them for our proofs.
Definition A.1 (Subdifferential and minimal section).
For any convex function $g:\mathbb{R}^{d}\to\mathbb{R}$, its subdifferential evaluated at $x$ is the set
@@
Recalling the definition of Moreau-Yosida approximation in Definition 2.2 we have that $g^{\lambda}(y_{0})=g(y_{1})+\|y_{0}-y_{1}\|^{2}/(2\lambda)$; plugging this into the equation above gives
@@
Finally, using Salim et al. [2019, Lemma 9] which states that $g^{\lambda}(x)\geq g(x)-\lambda\|\nabla^{0}g(x)\|/2$, where $\nabla^{0}g$ is the minimal section introduced in Definition A.1, and combining  (33) and (34) we have
@@
On the other hand, by Lemmas A.5 and A.7 we have

### v3

@@ 5.2 Bayesian Neural Network
MYPGD $\mathbf{1.50\pm 0.77}$ $-1.02\pm 0.15$ $20$ $2.02$
MYIPLA $2.00\pm 0.85$ $-1.07\pm 0.19$ $22$ $2.27$
PIPGLA $2.00\pm 0.75$ $\mathbf{-0.96\pm 0.09}$ $33$ $\mathbf{1.73}$
PGD $2.00\pm 0.98$ $-0.98\pm 0.10$ $\mathbf{19}$ $8.80$
SOUL $6.37\pm 1.56$ $-3.50\pm 2.43$ $5${55}55 $12.09$
IPLA $1.99\pm 1.01$ $-1.01\pm 0.25$ $\mathbf{19}$ $11.70$
5.3 Image Deblurring
The objective of image deconvolution is to recover a high-quality image from a blurred and noisy observation $y=Hx+\varepsilon$, where $H$ is a circulant blurring matrix and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$. This inverse problem is ill-conditioned, a challenge that Bayesian methods address by incorporating prior knowledge. A common choice is the total variation prior, which promotes smoothness and is defined as $TV(x)=\|\nabla_{d}x\|_{1}$, where $\|\cdot\|_{1}$ is the $\ell_{1}$ norm and $\nabla_{d}$ is the two-dimensional discrete gradient operator. However, the strength of this prior depends on a hyperparameter $\theta$ that typically requires manual tuning. Instead of fixing this parameter manually, we estimate its optimal value. Thus, the posterior distribution for the model takes the form $p_{\theta}(x|y)\propto C(\theta)\exp\left(-\|y-Hx\|^{2}/(2\sigma^{2})-e^{\theta}TV(x)\right)$. For the experiments, we use the algorithm proposed by [Douglas and Rachford, 1956] to numerically evaluate the proximal operator of the total variation norm. Qualitative results are presented in Figure 3, with additional results provided in Appendix E.4.
(a) Original
(b) Blurred
(c) MYIPLA
Figure 3: Image deblurring experiment.
6 Conclusion
Our algorithms present a novel approach for handling Bayesian models arising from different types of non-differentiable regularisations, including Lasso, elastic net, nuclear-norm and total variation norm. While our theoretical guarantees are established under strong convexity assumptions, in practice, our methods perform well under more general conditions and demonstrate robustness and stability across a range of regularisation parameter values. Moreover, unlike standard Langevin algorithms—which fail to converge for light-tailed distributions [Roberts and Tweedie, 1996a]—our proximal variants remain effective, due to the implicit regularisation introduced by the proximal map. Future work holds many promising avenues. Our theoretical framework can be extended to the non-convex setting using recent non-convex optimisation bounds [Zhang et al., 2023] and their non-differentiable adaptations. Additionally, our novel bounds on the difference between the true minimiser and the minimiser of the Moreau-Yosida approximation can also be used within recent multiscale approaches to extend them to non-differentiable settings, see, e.g., Akyildiz et al. [2024].
Acknowledgements.
@@
PCE is supported by EPSRC through the Modern Statistics and Statistical Machine Learning (StatML) CDT programme, grant no. EP/S023151/1. FRC gratefully acknowledges the “de Castro" Statistics Initative at the Collegio Carlo Alberto and the Fondazione Franca e Diego de Castro.
@@ A.1 Approximation of minimisers
Before proceeding with the proof of Theorem 4.1 and 4.2, we derive a result controlling the distance between the maximiser of $p_{\theta}(y)=\int_{\mathbb{R}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\star}$, and the maximiser of $p_{\theta}^{\lambda}(y)=\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\lambda,\star}$.
@@
$k_{\lambda}(\theta):=\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x,\qquad\qquad k(\theta):=\int_{\mathbb{R}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$
@@
i.e. positive definite due to Remark 3 and assumption A5. Thus the Jacobian of $\phi$ w.r.t. $w$ is invertible. Hence, the implicit function theorem shows that there is some locally defined $\zeta$, such that $\zeta(0)=v$ and $\phi(\lambda,\zeta(\lambda))=0$. Furthermore,
@@ A.2 MYIPLA
The proof follows from Karatzas and Shreve [1991] and Akyildiz et al. [2025, Proposition 1]. ∎
@@ A.3 PIPULA
for every $x,y\in\mathbb{R}^{d}$. By plugging this result into the proof of Akyildiz et al. [2025, Lemma B.1] we obtain an equivalent result to that of Akyildiz et al. [2025, Proposition 5]:
@@ A.4 PIPGLA
The concentration term $W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})$ is analysed in Lemma A.14 and Theorem A.15. For the term $W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$, we derive a novel bound. The roadmap for obtaining this latter bound is as follows:
  1. 1.
    We first analyse the PGLA updates targetting a distribution $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}^{\lambda}))$ in $\mathbb{R}^{d}$ given by
    $Z_{n}$ $=X_{n}-\gamma\nabla g_{1}(X_{n}),$
    $Y_{n}$ $=Z_{n}+\sqrt{2\gamma}\;\xi_{n},$
    $X_{n+1}$ $=\text{prox}_{g_{2}}^{\lambda}(Y_{n}).$
    We provide an error on $W_{2}(\mathcal{L}({X_{n+1}}),\pi)$ where $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}))$ for arbitrary $\lambda$ and $\gamma$. One key idea for this is the use of the minimal section which quantifies the least norm element in the subdifferential set $\partial g(x)$ introduced in Definition 3.
  2. 2.
    We show in Corollary A.12 that the results established in 1. can be applied to a proximal gradient scheme in which the noise is scaled by $\sqrt{N}$
    $V_{n+1/2}$ $=V_{n}-\gamma\nabla g_{1}(V_{n})+\sqrt{\frac{2\gamma}{N}}\;\xi_{n},$
    $V_{n+1}$ $=\text{prox}_{g_{2}}^{\lambda}(V_{n+1/2})=V_{n+1/2}-\lambda\nabla g_{2}^{\lambda}(V_{n+1/2}).$
  3. 3.
    $Z_{n+1/2}$ $=Z_{n}-\gamma\nabla G_{1}(Z_{n})+\sqrt{\frac{2\gamma}{N}}\xi_{n+1},$
    $Z_{n+1}$ $=\text{prox}_{G_{2}}^{\lambda}(Z_{n+1/2})=Z_{n+1/2}-\lambda\nabla G_{2}^{\lambda}(Z_{n+1/2}).$
    where $G_{1}$ and $G_{2}$ are defined as
    Since $G_{1}$ and $G_{2}^{\lambda}$ preserve the properties (strong convexity and Lipschitzness) of $g_{1}$ and $g_{2}^{\lambda}$ (see Corollary A.13), we use the result in 2. to bound the error $W_{2}(\mathcal{L}({Z_{n}}),\pi^{N})$. Finally, using a data processing inequality it follows $W_{2}(\mathcal{L}({\theta_{n}^{N}}),\pi_{\Theta}^{N})\leq W_{2}(\mathcal{L}({Z_{n}}),\pi^{N})$.
@@ A.4.1 Error bound for proximal gradient Langevin algorithm
For convenience we consider the following decomposition of the PGLA update targeting a distribution $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}^{\lambda}))$ over $\mathbb{R}^{d}$
@@
We are going to derive a bound for $W_{2}(\mathcal{L}({X_{n+1}}),\pi)$, where $\mathcal{L}({X_{n+1}})$ denotes the distribution of the random variable $X_{n+1}$. For every $\pi$-integrable function $g:\mathbb{R}^{d}\to\mathbb{R}$, we define $\mathcal{E}_{g}(\pi)=\int g\mathrm{d}\pi$ and we denote $\mathcal{F}=\mathcal{E}_{g_{1}+g_{2}}+\mathcal{H}$, where $\mathcal{H}$ is the negative entropy $\mathcal{H}(\pi)=\int\log\pi\;\mathrm{d}\pi$. We also introduce the subdifferential of a convex function and its minimal section, since we will use them for our proofs.
Definition 3.
[Subdifferential and minimal section] For any convex function $g:\mathbb{R}^{d}\to\mathbb{R}$, its subdifferential evaluated at $x$ is the set
@@
Recalling the definition of Moreau-Yosida approximation in Definition 2 we have that $g^{\lambda}(y_{0})=g(y_{1})+\|y_{0}-y_{1}\|^{2}/(2\lambda)$; plugging this into the equation above gives
@@
Finally, using Salim et al. [2019, Lemma 9] which states that $g^{\lambda}(x)\geq g(x)-\lambda\|\nabla^{0}g(x)\|/2$, where $\nabla^{0}g$ is the minimal section introduced in Definition 3, and combining  (30) and (31) we have
@@
On the other hand, by Lemmas A.7 and A.9 we have